nlknnkjklm:

After completing the fraction division 8 divided by StartFraction 1 Over 6 EndFraction, Patel found a quotient of mc025-2.jpg. He then used the multiplication below to check his work. StartFraction 1 Over 48 EndFraction times 6 = StartFraction 1 Over 48 EndFraction times StartFraction 6 Over 1 EndFraction = StartFraction 6 Over 48 EndFraction = StartFraction 1 Over 8 EndFraction Which is the most accurate description of Patel’s work?

snowflake0531:

Can you take a screenshot of your question?

mhanifa:

I found this

1 attachment
mhanifa:

Patel should have performed division: 8 : 1/6 = 8*6 = 48 And the verification should be like this: 48 * 1/6 = 8 What do you think is the correct answer choice?